# Explainable Artificial Intelligence for Autonomous Driving: A Comprehensive Overview and Field Guide for Future Research Directions

### v4

@@ 5.3 Vision-based explanations
whether erroneous actions are due to incorrect velocity or calculation of direction. Finally, Cost Map visualization describes the traffic scene via a top-down view. The architecture is evaluated on a large real-driving dataset consisting of 6,500 traffic scenarios with 1.4 million frames and collected across several cities in North America, and measuring traffic rule violation, closeness to human trajectory, and collision. The authors also carry out an ablation study and show the impact of different overrides, input horizons, and training losses on end-to-end learning.
@@
Xu et al. [125] propose object-induced actions with explanations for predictions of an autonomous car. The authors introduce a new dataset called BDD-OIA, as an extension of the BDD100K dataset [147]; this extension is annotated with 21 explanation templates on a set of 4 actions. Their multi-task formulation for predicting actions also improves the accuracy of action selection. The CNN architecture further unifies
Figure 9: Human advice to a vehicle for appropriate action. Source: [126].
reasoning on action-inducing objects and the context of scenes globally. The empirical results of the study on the introduced BDD-OIA dataset show that the explainability of the architecture also enhances action-inducing object recognition, resulting in better self-driving.
@@
As a more recent vision-to-text approach, Atakishiyev et al. [144] employ the visual question answering (VQA) mechanism to explain autonomous driving actions. They train an RL agent and generate driving data showing the self-driving car’s motion from its field of view. They further convert this video to image sequences, manually annotate the images with question-answer (QA) pairs, and encode questions and images with LSTM and pre-trained VGG-19 [148], respectively. The experimental results on five action categories show that VQA is a straightforward, effective, and human-interpretable approach to justify autonomous driving actions. Leveraging frontal images for interpretable decision-making has further been explored by subsequent studies as well [27, 143, 146].
@@
While the mentioned studies focus on vision-based explanations of already obtained predictions of the model, there have been some recent studies paying attention to counterfactual explanations. In the context of automated driving, counterfactual analysis can be described with such an exemplary question: “Given the driving scene, how can it be modified so that the vehicle keeps driving instead of stopping ?” In other words, given the input, counterfactual analysis intends to figure out the distinguished features in this input that cause the model to make a certain prediction by envisioning modification of those features would cause the model to make a different prediction (e.g., Figure 10). Thus, in this case, the predictions obtained by the existing model and the imagined model become contrastive. As the application of counterfactual intervention, Li et al. [127] presents an approach to find out risk objects that result in particular driving behavior. Their method, formalized as a Functional Causal Model (FCM), shows that the random elimination of some objects from the scene changes the driving decision to the contrastive prediction, such as from the “Stop” to “Go” command. In further work, Jacob et al. [136] introduce the STEEX model that uses a pre-trained generative model to produce counterfactual rationales by modifying
the style of the scene while retaining the structure of the driving scene. Finally, as further enhancement of STEEX, Zemni et al. [149] propose a method called OCTET that generates object-aware counterfactual explanations without depending on the structural layout of the driving scene as backpropagation can optimize the spatial positions of the provided instances.
@@
Overall, we observe a significant focus on perception-based explanations of autonomous driving systems, as such explanations provide an opportunity to better understand how accurately a self-driving vehicle senses the operational environment. We show a summary of vision-based explanations for AVs in Table 2.
5.4 RL and IL-based explanations
@@
Explaining how perceived environmental states are mapped to actions has also recently received attention in the autonomous driving community. In this regard, the field of explainable RL (XRL) is a relatively new and emerging research avenue on XAI [150, 151, 152]. Like vision-based explanations, XRL techniques also aim to provide some forms of justifications on chosen actions of a vehicle either via intrinsically interpretable design or post-hoc explanations. One of the early works in this context is the Semantic Predictive Control (SPC) framework [153], where the authors propose a data-efficient policy learning approach that predicts future semantic segmentation and provides visual explanations of a learned policy. The framework concatenates multi-scale intermediate features from RGB with tiled actions. The joined modules are then fed into the multi-scale prediction model that predicts future features. Finally, in the last part of the pipeline, the information prediction module inputs the latent feature representation and outputs driving signals alongside the semantic segmentation of the scene.
@@
Chen et al. [154] introduce a sequential latent environment model learned with RL and a probabilistic graphical model-based approach that can interpret autonomous cars’ actions via a bird-eye mask. They use video cameras and LIDAR images as input in the CARLA simulator. For the purpose of interpretability of actions and explainability of a learned policy, they generate a bird-eye mask (i.e., Figure 11). Their model outperforms the used baseline models - DQN, DDPG, TD3, and SAC. Similarly, Wang et al. [155] propose an interpretable end-to-end vision-based motion planning (IVMP) to interpret the underlying actions of a self-driving vehicle. They use semantic maps of a bird-view space in order to plan the motion trajectories of an autonomous car. Moreover, the IVMP approach uses an optical flow distillation network that can improve the real-time performance of the network. The experiments conducted on the nuScenes dataset [156] show the superiority of the proposal over modern approaches in semantic map segmentation and imitation of human drivers. In another probabilistic decision-making model, Wang et al. [157] approach lane merging task as a dynamic process and integrate internal states into joint Hidden Markov Model (HMM) and Gaussian Mixture Regression (GMM). The experiments collected on the INTERACTION dataset [158] demonstrate the efficiency of the proposed technique and show that merging at highway on-ramps can be delineated by three interpretable internal states in terms of the absolute speed of a vehicle while merging.
Rjoub et al. [159] have shown that federated deep RL combined with XAI can lead to trusted autonomous
Table 3: Explainable RL and IL-based studies for AVs
@@
driving. They use a federated learning approach for decision-making and leverage edge computing that enables different devices to train an ML model in a collaborative manner. The model is first developed on the parameter server and further broadcasted to other devices. Then, global ML methods intake updates from these devices regularly, and the process continues until the model performs well enough on driving tasks. Yang et al. [169] have also shown that an intrinsically interpretable RL agent for autonomous driving can be achieved via reward consistency aiming to resolve the gradient disconnection in reward-action mapping.
Finally, within the exploration via interaction context, a few studies have employed various forms of IL techniques for explainable autonomous driving. Cultrera et al. [161] present conditional imitation learning with an end-to-end visual attention model, which identifies those parts of images that have a higher influence on predictions. They test their architecture on the CARLA simulator [172] on four tasks - go straight, turn left, turn right, and follow the lane. Their ablation study focused on box type importance and fixed grid analysis to get an attention map on the images shows that integrated imitation learning and
Table 4: Studies on decision tree-based explanations for AVs
@@
attention model enables a car to drive safely and perform relevant maneuvers in real-time.
Leveraging vision for interpretability of an agent’s actions, Teng et al. [177] propose hierarchical interpretable IL (HIIL) technique that unifies bird eye view (BEV) mask with the steering angle to perform actions in complex situations as an end-to-end autonomous driving pipeline. They construct their method as a two-phase task: In the first phase, the pre-trained BEV model is used to interpret the driving environment. Then imitation learning takes the latent features of BEV mask from the first phase and combines them with a steering angle acquired through the Pure-Pursuit algorithm. The experiment performed on the CARLA simulator shows that the proposed method enhances the interpretability and robustness of driving in various circumstances.
Moreover, Renz et al. [163] introduce PlanT, a rigorous IL-based approach that uses transformers for planning. PlanT is able to explain its action decision by recognizing the most important object in its driving segment and outperforms state-of-the-art work on CARLA’s Longest6 Benchmark by 10 points (See [178] for a visual demonstration). In a more recent work in this context, Liu et al. [171] show that combining Signal Temporal Logic with generative IL is also an effective approach for interpretable policy for autonomous driving. Overall, while building interpretable by-design RL or IL agent for autonomous driving is a challenging task, employing external methods, such as logic and vision, can help achieve explainability for the agent’s actions. A summary of RL and IL-based explanations is shown in Table LABEL:tab:rl_il_exp.
@@ 5.5 Decision tree-based explanations
Table 5: Studies on logic-based explanations for AVs
@@
5.6 Logic-based explanations
While the interpretability of a deployed autonomous driving control model has been the dominant direction for research, there have also been attempts to verify the safety of self-driving vehicles with logical reasoning. In this regard, Corso and Kochenderfer [182] present a technique to identify interpretable failures of autonomous cars. They use signal temporal logic expressions to describe failure cases of an autonomous car in unprotected left turn and pedestrian crossing scenarios. For this purpose, the authors use genetic programming to optimize signal temporal logic expressions that acquire disturbances trajectories causing a vehicle to fail in decisive action. The experimental results show that the proposed approach is effective in interpreting the safety validation of a car.
5.7 User study-based explanations
Finally, except for practical applications, some investigations involve users in case studies to understand the effective strategies for explanation generation in autonomous driving tasks. The key idea of a user study is that getting people’s input in designated driving tasks can help improve the adequacy and quality of explanations in autonomous driving. Wiegand et al. [187] perform a user study that identifies a mental model of users for determining an effective practical implementation of an explanation interface. The main research question here is to understand what components need to be visualized in a vehicle so the user can comprehend the decisions of self-driving vehicles. The study discloses that combining an expert mental model with a user mental model as a target mental model enhances the situation awareness of the drivers. Furthermore, Wiegand et al. [188] investigate situations, in which explanations are needed and methods pertinent to these situations. They spot seventeen scenarios where a self-driving vehicle behaves unexpectedly. Twenty-six participants are selected to validate these situations in the CarMaker driving simulator to provide insights into drivers’ need for explanations. As a result of the user study, the authors identify six groups to highlight the primary concerns of drivers with these unexpected behaviors, namely emotion and evaluation, interpretation and reason, the capability of a self-driving car, interaction, driving forecasting, and request times for explanations.
@@ 5.7 User study-based explanations
Wang et al. [130] propose an approach that enables a human driver to provide scene forecasting to an intelligent driving system using a purposeful gaze. They develop a graphical user interface to understand the effect of human drivers on the prediction and control of an intelligent vehicle. A simulator is used to
Table 6: Studies on user study-based explanations for AVs
@@
test and verify three driving situations where a human driver’s input can improve safety of self-driving.
While the preliminary studies and further works on explainable autonomous driving primarily focused on a combination of various AI techniques revisited above, large language models (LLMs) and vision-language models (VLMs) have recently emerged as a novel paradigm to interpreting AVs decisions and describing traffic scenes. Built on top of Foundation Models, such as GPT [193] and BERT [194], there has been significant progress on building both domain-agnostic and domain-specific LLMs (e.g., GPT-3 [195], GPT-4 [196], LLAMA [197], LLAMA-2 [198], Vicuna [199], Alpaca [200], Claude [201]) and VLMs (e.g., Flamingo [202], LLaVa [203], PaLM-E [204], Video-LLaVA [205], Video-LLaMA [206], Gemini [207], Claude 3 [208]). In this sense, there have been tremendous efforts to build language and vision-language models on top of these base models for interpretable autonomous driving. Overall, based on the recent trend in leveraging these large models for the interpretability purpose in autonomous driving, we observe the following directions:
Presenting live natural language explanations during the trip: The promising work in this context is Wayve’s LINGO-1 [209] architecture. LINGO-1 is a vision-language-action (VLAM) model that provides live natural language explanations for describing a vehicle’s chosen actions in end-to-end autonomous driving. Trained on diverse multimodal (vision and language) datasets, LINGO-1 can describe action decisions and causal factors inducing these actions. The advantage of LINGO-1 architecture is that its explanations are concise, informative, and reflect temporal changes in the driving environment. An example of LINGO-1’s live and temporal explanations is shown in Figure 12. The Wayve team has also achieved live linguistic explanations for autonomous driving in a simulation environment [210].
Video Question Answering as a reasoning technique: An essential characteristic of modern AVs is the
Table 7: Large language and vision-language based explanations for AVs

### v5

@@ 4.3 Visual explanations
Zeng et al.’s [70] architecture learns to drive an autonomous vehicle safely by following traffic rules, including interaction with road users, yielding, and traffic signals. They use raw LIDAR data and an HD map to generate interpretable representations as 3D detection of objects, anticipated future trajectories, and cost map visualizations. 3D detection instances provide descriptive information so that the model understands the operational environment. Motion forecasting, measured as L1 and L2 distances, explains whether erroneous actions are due to incorrect velocity or calculation of direction. Finally, Cost Map visualization describes the traffic scene via a top-down view. The architecture is evaluated on a large real-driving dataset consisting of 6,500 traffic scenarios with 1.4 million frames and collected across several cities in North America, and measuring traffic rule violation, closeness to human trajectory, and collision. The authors also carry out an ablation study and show the impact of different overrides, input horizons, and training losses on end-to-end learning.
@@
Xu et al. [72] propose object-induced actions with explanations for predictions of an autonomous car. The authors introduce a new dataset called BDD-OIA, as an extension of the BDD100K dataset [107]; this extension is annotated with 21 explanation templates on a set of 4 actions. Their multi-task formulation for predicting actions also improves the accuracy of action selection. The CNN architecture further unifies reasoning on action-inducing objects and the context of scenes globally. The empirical results of the study on the introduced BDD-OIA dataset show that the explainability of the architecture also enhances action-inducing object recognition, resulting in better self-driving.
@@
As a more recent vision-to-text approach, Atakishiyev et al. [90] employ the visual question answering (VQA) mechanism to explain autonomous driving actions. They train an RL agent and generate driving data showing the self-driving car’s motion from its field of view. They further convert this video to image sequences, manually annotate the images with question-answer (QA) pairs, and encode questions and images with LSTM [108] and pre-trained VGG-19 [109], respectively. The experimental results on five action categories show that VQA is a straightforward, effective, and human-interpretable approach to justify autonomous driving actions. Leveraging frontal images for interpretable decision- making has further been explored by subsequent studies as well [18, 8, 91].
@@
While the mentioned studies focus on vision-based explanations of already obtained predictions of the model, there have been some recent studies paying attention to counterfactual explanations. In the context of automated driving, counterfactual analysis can be described with such an exemplary question: “Given the driving scene, how can it be modified so that the vehicle keeps driving instead of stopping ?” In other words, given the input, counterfactual analysis intends to figure out the distinguished features in this input that cause the model to make a certain prediction by envisioning modification of those features would cause the model to make a different prediction (e.g., Figure 4). Thus, in this case, the predictions obtained by the existing model and the imagined model become contrastive. As the application of counterfactual intervention, Li et al. [74] presents an approach to find out risk objects that result in particular driving behavior. Their method, formalized as a Functional Causal Model (FCM), shows that the random elimination of some objects from the scene changes the driving decision to the contrastive prediction, such as from the “Stop” to “Go” command. In further work, Jacob et al. [83] introduce the STEEX model that uses a pre-trained generative model to produce counterfactual rationales by modifying the style of the scene while retaining the structure of the driving scene. Finally, as further enhancement of STEEX, Zemni et al. [86] propose a method called OCTET that generates object-aware counterfactual explanations without depending on the structural layout of the driving scene as backpropagation can optimize the spatial positions of the provided instances.
@@
Overall, we observe a significant focus on perception-based explanations of autonomous driving systems, as such explanations provide an opportunity to better understand how accurately a self-driving vehicle senses the operational environment. We show a
Figure 5: Human advice to the car for relevant action. Source: [73].
summary of vision-based explanations for AVs in Table I.
4.4 Reinforcement learning and Imitation learning-based explanations
@@
Explaining how perceived environmental states are mapped to actions has also recently received attention in the autonomous driving community. In this regard, the field of explainable reinforcement learning (XRL) is a relatively new and emerging research avenue on XAI [110, 111, 112]. Like vision-based explanations, XRL techniques also aim to provide some forms of justifications on chosen actions of a vehicle either via intrinsically interpretable design or post-hoc explanations. One of the early works in this context is the Semantic Predictive Control (SPC) framework [113], where the authors propose a data-efficient policy learning approach that predicts future semantic segmentation and provides visual explanations of a learned policy. The framework concatenates multi-scale intermediate features from RGB with tiled actions. The joined modules are then fed into the multi-scale prediction model that predicts future features. Finally, in the last part of the pipeline, the information prediction module inputs the latent feature representation and outputs driving signals alongside the semantic segmentation of the scene.
@@
Chen et al. [114] introduce a sequential latent environment model learned with RL and a probabilistic graphical model-based approach interpreting autonomous cars’ actions via a bird-eye mask. They use video cameras and LIDAR images as input in the CARLA simulator [115]. For the purpose of interpretability of actions and explainability of a learned policy, they generate a bird-eye mask (i.e., Figure 6). Their model outperforms the used baseline models - DQN, DDPG, TD3, and SAC. Similarly, Wang et al. [116] propose an interpretable end-to-end vision-based motion planning (IVMP) to interpret the underlying actions of a self-driving vehicle. They use semantic maps of a bird-view space in order to plan the motion trajectories of an autonomous car. Moreover, the IVMP approach uses an optical flow distillation network that can improve the real-time performance of the network. The experiments conducted on the nuScenes dataset [117] show the superiority of the proposal over modern approaches in semantic map segmentation and imitation of human drivers. In another probabilistic decision-making model, Wang et al. [118] approach lane merging task as a dynamic process and integrate internal states into joint Hidden Markov Model (HMM) and Gaussian Mixture Regression (GMM). The experiments conducted on the
TABLE II: Studies on reinforcement learning and imitation learning-based explanations for AVs
@@
INTERACTION dataset [132] demonstrate the efficiency of the proposed technique and show that merging at highway on-ramps can be delineated by three interpretable internal states in terms of the absolute speed of a vehicle while merging.
Rjoub et al. [122] have shown that federated deep RL combined with XAI can lead to trusted autonomous driving. They use a federated learning approach for decision-making and leverage edge computing that enables different devices to train an ML model in a collaborative manner. The model is first developed on the parameter server and further broadcasted to other devices. Then, global ML methods intake updates from these devices regularly, and the process continues until the model performs well enough on driving tasks. Yang et al. [129] have also shown that an intrinsically interpretable RL agent for autonomous driving can be achieved via reward consistency aiming to resolve the gradient disconnection in reward-action mapping.
Finally, within the exploration via interaction context, a few studies have employed various forms of imitation learning (IL) techniques for explainable autonomous driving. Cultrera et al. [120] present conditional imitation learning with an end-to-end visual attention model, which identifies those parts of images that have a higher influence on predictions. They test their architecture in the CARLA simulator on four tasks - go straight, turn left, turn right, and follow the lane. Their ablation study focused on box type importance and fixed grid analysis to get an attention map on the images shows that integrated imitation learning and attention model enables a car to drive safely and perform relevant maneuvers in real-time.
Leveraging vision for interpretability of an agent’s actions, Teng et al. [133] propose hierarchical interpretable IL (HIIL) technique
@@
that unifies bird eye view (BEV) mask with the steering angle to perform actions in complex situations as an end-to-end autonomous driving pipeline. They construct their method as a two-phase task: In the first phase, the pre-trained BEV model is used to interpret the driving environment. Then imitation learning takes the latent features of BEV mask from the first phase and combines them with a steering angle acquired through the Pure-Pursuit algorithm. The experiment performed in the CARLA simulator shows that the proposed method enhances the interpretability and robustness of driving in various circumstances.
Moreover, Renz et al. [123] introduce PlanT, a rigorous IL-based approach that uses transformers for planning. PlanT is able to explain its action decision by recognizing the most important object in its driving segment and outperforms state-of-the-art work in CARLA’s Longest6 Benchmark by 10 points (See [134] for a
TABLE III: Studies on decision tree-based explanations for AVs
@@
visual demonstration). In a more recent work in this context, Liu et al. [131] show that combining Signal Temporal Logic with generative IL is also an effective approach for interpretable policy for autonomous driving. Overall, while building interpretable by-design RL or IL agent for autonomous driving is a challenging task, employing external methods, such as logic and vision, can help achieve explainability for the agent’s actions. A summary of RL and IL-based explanations is shown in Table LABEL:tab:rl_il_exp.
@@ 4.5 Decision tree-based explanations
4.6 Logic-based explanations
While the interpretability of a deployed autonomous driving control model has been the dominant direction for research, there have also been attempts to verify the safety of self-driving vehicles with logical reasoning. In this regard, Corso and Kochenderfer [141] present a technique to identify interpretable failures of autonomous cars. They use signal temporal logic expressions to describe failure cases of an autonomous car in unprotected left turn and pedestrian crossing scenarios. For this purpose, the authors use genetic programming to optimize signal temporal logic expressions that acquire disturbances trajectories causing a vehicle to fail in its decisive action. The experimental results show that the proposed approach is effective in interpreting the safety validation of a car.
4.7 User study-based explanations
Some investigations involve users in case studies to understand the effective strategies for explanation generation in autonomous driving tasks. The key idea of a user study is that getting people’s input in designated driving tasks can help improve the adequacy and quality of explanations in autonomous driving. Wiegand et al. [146] perform a user study that identifies a mental model of users for determining an effective practical implementation of an explanation interface. The main research question here is to understand what components need to be visualized in a vehicle so the user can comprehend the decisions of self-driving vehicles. The study discloses that combining an expert mental model with a user mental model as a target mental model enhances the situation awareness of the drivers. Furthermore, Wiegand et al. [147] investigate situations, in which explanations are needed and methods pertinent to these situations. They spot seventeen scenarios where a self-driving vehicle behaves unexpectedly. Twenty-six participants are selected to validate these situations in the CarMaker driving simulator to provide insights into drivers’ need for explanations. As a result of the user study, the authors identify six groups to highlight the primary concerns of drivers
TABLE IV: Studies on logic-based explanations for AVs
@@ 4.7 User study-based explanations
with these unexpected behaviors, namely emotion and evaluation, interpretation and reason, the capability of a self-driving car, interaction, driving forecasting, and request times for explanations.
@@
Wang et al. [77] propose an approach that enables a human driver to provide scene forecasting to an intelligent driving system using a purposeful gaze. They develop a graphical user interface to understand the effect of human drivers on the prediction and control of an intelligent vehicle. A simulator is used to test and verify three driving situations where a human driver’s input can improve safety of self-driving.
Finally, while the preliminary studies and further works on explainable autonomous driving primarily focused on a combination of various AI techniques revisited above, large language models (LLMs) and vision-language models (VLMs) have recently emerged as a novel paradigm to interpreting AVs decisions and describing traffic scenes. Built on top of Foundation Models, such as GPT [152] and BERT [153], there has been significant progress on building both domain-agnostic and domain-specific LLMs (e.g., GPT-3 [154], GPT-4 [155], LLAMA [156], LLAMA-2 [157], Vicuna [158], Alpaca [159], Claude [160]) and VLMs (e.g., Flamingo [161], LLaVa [162], PaLM-E [163], Video-LLaVA [164], Video-LLaMA [165], Gemini [166], Claude 3 [167]). In this sense, there have been tremendous efforts to build language and vision-language models on top of these base models for interpretable autonomous driving. Overall, based on the recent trend in leveraging these large models for the interpretability purpose in autonomous driving, we observe the following directions:
Presenting live natural language explanations during the trip: The promising work in this context is Wayve’s LINGO-1 [168] and LINGO-2 [169] architectures. LINGO-1 is a vision-language-action (VLAM) model that provides live natural language explanations for describing a vehicle’s chosen actions in end-to-end autonomous driving. Trained on diverse multimodal (vision and language) datasets, LINGO-1 can describe action decisions and causal factors inducing these actions. The advantage of LINGO-1 architecture is that its explanations are concise, informative, and reflect temporal changes in the driving environment. The Wayve team has also achieved live linguistic explanations for autonomous driving in a simulation environment [170].
Video Question Answering as a reasoning technique: An essential characteristic of modern AVs is the consideration of human factors in the design and development of this technology and having effective human-machine alignment for trustworthy autonomous driving. In this sense, it is crucial that users-in-the-loop have some form of interaction with AVs during a journey. Motivated by this concept, some recent works have approached conversational user interface between people on board and AVs as a Video Question Answering (VideoQA) task [171, 172, 173]. Asking questions about the behavior of an autonomous system is a part of our intuition, and in the context of AVs, getting answers to traffic-related situations and autonomous car action-related questions can help users have a comfortable and reliable journey. Other practical applications of LLMs and VLMs are interpretable motion planning [174], chain-of-thought-based reasoning for control and decision-making [175, 176], action justification with a control signal [177], predicting the intent of other traffic actors via visual reasoning and cues [178], and understanding the role of video transformer based-explanations on safety of autonomous driving [179] (see Table LABEL:tab:llm_vlm_exp for more details on LLM and VLM-based explanations).
Overall, as an emerging AI technology, LLMs and VLMs have tremendously benefitted AVs from the interpretability aspect, as described in the above studies. However, it is also worthwhile to mention that there are still spaces for improvement of these models as fictitiously generated explanations (e.g., hallucinations) may have serious safety implications and high-stakes consequences for self-driving actions and human life. We describe such caveats and potential solutions in Section 6 as future work.
A high-level overview of all these studies indicates driving explanations are generally multi-modal, context-dependent, and task-specific. Moreover, end-to-end learning has become even more popular for highly autonomous decision-making owing to the combination of powerful deep-learning approaches and overall safety benefits. Based on the insights from the reviewed literature, we can define explainable autonomous driving as follows:
TABLE V: User study-based explanations for AVs
@@ 4.8 Large Language Models and Vision-Language Models-based explanations
Driven by this definition and the state-of-the-art works in the above sections, we present a general and conceptual XAI framework for end-to-end autonomous driving aligned with industrial trends and show the necessary components and process steps to achieve regulatory-compliant AVs in the next generation.
5 An XAI framework: integrating end-to-end control, safety, and explanations
We present a general framework in which methods for developing XAI, end-to-end learning, and safety components are combined to inform processes of regulatory principles. Each of these components has a concrete role in our framework. In our recent study [181], we have covered a brief description of end-to-end learning for AVs. We extend the scope of that work and describe the essential elements of end-to-end autonomous driving, and the role of and potential challenges with explanations in such a setting. We describe these individual components as follows:
$E=\{e_{1},e_{2},...e_{n}\},$
and a compendium of actions
$A=\{a_{1},a_{2},...a_{n}\},$
$C:E\mapsto A.$
$e\in E$
to a relevant action
$a\in A.$
$srC=f(eeC,A).$
This requirement could be as pragmatic as some inspection of individual vehicle safety (for example, verifying basic safety functions of an individual vehicle for re-licensing). That said, this concept should be considered as a thorough compliance testing of $eeC$ components from vehicle manufacturers to certify their public safety under international and/or national transportation guidelines such as [32] and [37]. The general principles for acceptable functional safety of road vehicles are defined by the ISO 26262 standard [43]. According to this standard, there should be a safety certification development with evidence-based rationales: the vehicle should be able to meet the established functional safety requirement in its operational context. Part 6 of the ISO 26262 standard [182] is dedicated to end-product development for automotive applications within the software level. This guideline includes the design, development, testing, and verification of software systems in automotive applications. Based on these standards, there seem to be two fundamental approaches to confirming regulatory compliance, which we label confirmation of compliance by “simulation,” and confirmation of compliance by “verification,”. These steps are aligned with our observation regarding the role of XAI in confirming regulatory compliance. In the case of the process of establishing regulatory compliance by simulation, the idea is that a selected set of autonomous actions can be simulated, and then assessed to be satisfactory. This approach is perhaps the most familiar, as it arises naturally from an engineering development trajectory, where the accuracy of simulators determines the quality of compliance (e.g., [183]). The confidence of the established compliance is a function of the accuracy and coverage of the simulation. However, this compliance process can be very expensive and prone to safety gaps, especially when consensus on the properties and scope of a simulation is difficult to achieve. Thus, in general, the simulation part can be considered a ”driving school” for AVs: The designed and developed learning software system should be tested rigorously
TABLE VI: Studies on large language models and vision-language models-based explanations for AVs